# Enhancing Hemodynamic Parameter Estimations: Nonlinear Blood Behavior in 4D Flow MRI

### v2

@@ 5 Discussion
Experiment 1 tested the necessity of a non-Newtonian model when computing key parameters such as those defined in (6). The significance of both experiments lies in assessing whether the Newtonian linear model ($\mu_{\mathrm{NF}}$) could serve as a surrogate for the power-law constitutive model, thus eliminating the need to account for nonlinear rheology. However, the differences observed in Figure 7 indicate that the two models are neither interchangeable, similar, nor equivalent. These differences reached up to 120% in some cases. These results suggests that, given the well-established shear-thinning behavior of blood in the literature, a constant viscosity model is insufficient for estimating hemodynamic parameters. The flow dynamics cannot be accurately characterized using a constant viscosity, underscoring the need to consider nonlinear models.
@@
Since we have established the necessity of accounting for the nonlinear behavior of blood, Experiment 2 evaluated the extent of the differences in WSS, $\dot{E}_{L}$, and OSI when using an incorrect (but usually employed), model—specifically, Newtonian (constant) viscosities ($\mu_{\mathrm{N}}$)—compared to the power-law viscosity model ($\mu_{\mathrm{PL}}$). Results showed significant differences between the estimations made with the two models (see Figures 9, 10, and 11), highlighting that using a constant viscosity model can lead to inaccurate predictions of hemodynamic parameters when nonlinear blood rheology is present. This was further confirmed in Figure 8, where WSS maps estimated from 4D Flow images demonstrated similar spatial distributions and Hct-level variations compared to their CFD simulation counterparts, though significant underestimations were observed at peak systole.
@@
However, the underestimation found at peak systole aligns with previous studies, which indicate that hemodynamic parameters, particularly WSS, are consistently underestimated when using 4D Flow MRI [68, 69, 66, 70, 71, 72, 64, 63]. Moreover, we observed the same behavior reported in [72] and [63]: WSS is significantly underestimated during systole (around 40% in our case), and these differences diminish during diastole. These findings provide confidence that all estimations made from images are faithful representations of the ground-truth estimations derived from CFD simulations.
@@
Having said that, it is important to note that this investigation focused solely on the use of the power-law model. Other viscosity models, such as the Carreau-like [13, 14, 15], Cross [16, 12], and Casson models [17], were not considered. However, as shown in the literature, the power-law model has provided similar results in terms of flow dynamics and hemodynamic parameters when compared to other models in vessels such as the coronary arteries and the aorta [73, 74]. It remains unclear whether it would be possible to differentiate between these models using 4D Flow MRI measurements due to the presence of partial volume effects, flow artifacts, and noise.
@@ 6 Conclusion
Our study highlights the importance of accounting for the nonlinear behavior of blood when estimating hemodynamic parameters from 4D Flow MRI using available CBC data. The conventional use of Newtonian or fitted-Newtonian viscosities is insufficient for capturing the complex dynamics of blood flow, particularly in major vessels like the aorta. This research shows that using the simple power-law model as the ground truth leads to significant differences in both simulations and synthetic 4D Flow MRI data compared to linear Newtonian (or fitted-Newtonian) models. These differences, across a wide range of Hct values, can be as high as 190% and 113% in WSS and $\dot{E}_{L}$ at systole, and as low as -72% and -74% at diastole, respectively. In in-vivo data, differences in WSS and $\dot{E}_{L}$ can reach up to -45% and -60% at systole, and range from -69% to 73% at diastole.
A key limitation is the lack of readily available power-law parameters for a wide range of Hct levels under physiological conditions. This highlights the need for future research to develop comprehensive databases of power-law parameters, as well as other rheological models, across different Hct values. Such databases would improve the accuracy of hemodynamic parameter estimations and contribute to the standardization of blood rheology research. In this regard, the methodology described in Section 2.4 is freely available at github.com/hmella/power-law-parameters.
Future work will focus on four main areas. First, we will explore the potential of blood viscosity maps as biomarkers by investigating the spatial distribution of viscosity in the cardiovascular system and its correlation with disease-related changes. Second, we will assess the impact of timely Hct measurements on the estimation of hemodynamic parameters. Third, we will conduct a statistical analysis across a wide range of Hct values to evaluate the robustness and clinical applicability of the proposed model in both volunteers and patients. Finally, we will investigate the use of other rheological models, such as Carreau-like, Cross, and Casson, for the estimation of hemodynamic parameters.
@@ 7 Acknowledgments
HM, JS, and EC acknowledge the financial support given by ANID-Fondecyt through the projects Iniciación en Investigación #11241098 and #11200481, and Regular #1210156, respectively. JS also acknowledges to ANID-Millennium Science Initiative Program ICN2021_004. FG acknowledges the funding PUCV Vinci DI-Iniciación 039.482/2024. We would like to kindly thank the Weierstraß-Institut für Angewandte Analysis und Stochastik (WIAS) for allowing the authors to run CFD simulations in their servers.
@@ Appendix A Hemodynamic parameters for each aortic segments
Figure 1: Mean values and standard deviations of WSS, $\dot{E}_{L}$, and OSI within each aortic segment, estimated from simulated 4D Flow images generated from CFD simulations using power-law ($\mu_{\mathrm{PL}}$) and fitted-Newtonian ($\mu_{\mathrm{NF}}$) viscosities. The estimations were performed using viscosity models and values consistent with those used in the CFD simulations that generated the synthetic 4D Flow images. Results are presented for both peak systole and diastole. Higher values and greater variability are observed in the ascending aorta due to the flow dynamics. The same differences shown in Figure 7 are clearly depicted here for each segment.

### v3

@@ 5 Discussion
It is noteworthy that for higher Hct values (i.e., higher viscosities), the error in the fitted power-law viscosities is greater (see RMSE values reported in Figure 2a). This is worth discussing because, for higher viscosities, there should be less experimental error in the measurements, implying greater certainty about the data. Moreover, this observation could suggest that the power-law model might not be the optimal choice to represent the experimental data. However, as mentioned in Section 2.2, this model was chosen based on its practicality: it can be fully defined using 4D Flow MR velocity measurements and a blood test, which is always available during the examination. Therefore, no additional experiments are required to determine the parameters of a more complex rheological model, nor is specialized equipment needed. This makes the power-law model suitable for adoption in any clinical context.
Also related to the fitting procedure, the adopted methodology for estimating Newtonian viscosities fitted by Hct yields a viscosity value of 0.97 mPa, which is lower than the plasma viscosity reported in the literature, typically ranging from 1 to 1.6 mPa [27, 28]. However, as mentioned earlier, the methodology employed in this study was the most appropriate choice given the available blood viscosity measurements across the Hct range considered, while also maintaining the physical consistency of the fit.
@@
Experiment 1 tested the necessity of a non-Newtonian model when computing key parameters such as those defined in (3.1) directly from synthetic velocity measurements obtained from synthetic 4D Flow MR images generated from CFD simulations. The significance of this experiment lies in assessing whether the Newtonian linear model ($\mu_{\mathrm{NF}}$) could serve as a surrogate for the power-law constitutive model, thereby eliminating the need to account for nonlinear rheology. However, the differences observed in Figure 8 indicate that the two models are neither interchangeable, similar, nor equivalent. In some cases, differences between models reached up to -63%. These results suggests that, given the well-established shear-thinning behavior of blood in the literature, a constant viscosity model is insufficient for estimating hemodynamic parameters from 4D Flow velocity measurements. Furthermore, the findings indicate that flow dynamics cannot be accurately characterized using a constant viscosity, highlighting the need to consider nonlinear models.
@@
Since we have established the necessity of accounting for the nonlinear behavior of blood, Experiment 2 evaluated the extent of the differences in WSS, $\dot{E}_{L}$, and OSI, estimated from velocity measurements obtained from synthetic 4D Flow MR images, when using an incorrect (but commonly employed) Newtonian (constant) viscosity ($\mu_{\mathrm{N}}$) compared to power-law viscosities ($\mu_{\mathrm{PL}}$) fitted to Hct values. Results showed significant differences between the estimations made with the two models (see Figures 10, 11, and 12), highlighting that using a constant viscosity model can lead to inaccurate predictions of hemodynamic parameters when nonlinear blood rheology is present. This was further supported by Figure 9, where WSS maps estimated from velocity measurements obtained from synthetic 4D Flow MR images demonstrated similar spatial distributions and Hct-level variations compared to their CFD simulation counterparts, although significant underestimations were observed at peak systole.
@@
However, the underestimation found at peak systole aligns with previous studies, which indicate that hemodynamic parameters, particularly WSS, are consistently underestimated when using 4D Flow MRI [87, 88, 85, 89, 90, 91, 83, 82]. Additionally, we observed the same behavior reported in [91] and [82]: WSS is significantly underestimated during systole (around 40% in our case), with these differences diminishing during diastole. These findings provide confidence that all estimations made from 4D Flow MR images are faithful representations of the ground-truth estimations derived from CFD simulations.
@@
It is worth mentioning that our results obtained from CFD simulations are consistent with findings reported in the existing literature. For instance, we observed that hemodynamic parameters increased with Hct in non-Newtonian simulations, which aligns with the results presented in [92] for carotid artery simulations. However, the same cannot be asserted for hemodynamic parameters estimated from 4D Flow MR images, as, to the authors’ knowledge, this is the first study to thoroughly assess this using both synthetic and in-vivo aortic 4D Flow MR data across a wide range of Hct values. However, the authors acknowledge the valuable contributions of 4D flow studies that incorporate non-Newtonian blood rheologies [47, 46].
@@
Having said that, it is important to note that this investigation focused solely on the use of the power-law model. Other viscosity models, such as the Carreau-like [19, 20, 21], Cross [22, 18], and Casson models [23], were not considered. However, as shown in the literature, the power-law model has produced similar results in terms of flow dynamics and hemodynamic parameters when compared to other models in vessels such as the coronary arteries and the aorta [93, 94], although it may not capture phenomena such as the plastic component of blood or inhomogeneous effects. That said, it remains unclear whether these differences between models could be detected using 4D Flow MRI measurements, given the presence of partial volume effects, flow artifacts, and noise.
@@ 6 Conclusion
Our study highlights the importance of accounting for the nonlinear behavior of blood when estimating hemodynamic parameters from 4D Flow MRI using available CBC data. The conventional use of Newtonian or Newtonian-fitted viscosities is insufficient for capturing the complex dynamics of blood flow, particularly in major vessels such as the aorta. This research demonstrates that using the simple power-law model as the ground truth leads to significant differences in both simulations and synthetic 4D Flow MRI data when compared to linear Newtonian (or Newtonian-fitted) models.
Since blood is intrinsically non-Newtonian, accounting for this behavior during the processing of 4D Flow MR images may enhance the accuracy and realism of hemodynamic parameter estimation. In this context, modeling viscosity using a power-law formulation emerges as a reasonable and straightforward alternative. Adopting this simple yet impactful methodological change could significantly improve the diagnosis and assessment of cardiovascular diseases through 4D Flow MR examinations, as health conditions involving variations in Hct would be directly reflected in the estimated hemodynamic parameters.
During the development of this study, a key challenge encountered was the lack of readily available power-law parameters for a wide range of Hct levels under physiological conditions. Although we highlight the existence of numerous notable studies assessing the role of Hct in blood rheology (see, for instance, [30, 25, 26, 27, 28]), they were not applicable to this research due to limitations such as a narrow Hct range, the use of different viscosity models, non-physiological measurement conditions, or the unavailability of raw data. This limitation highlights the need for future research aimed at developing comprehensive databases of power-law parameters, as well as those for other rheological models, across varying Hct values. Such resources would enhance the accuracy of hemodynamic parameter estimation and promote the standardization of blood rheology research. In this regard, the methodology described in Section 2.4 is freely available at www.github.com/hmella/power-law-parameters.
We also identified other limitations of this study. For instance, viscosity models such as the Carreau-like, Cross, and Casson were not considered in this investigation. Additionally, it remains unclear whether CBC exams conducted on a different day from the 4D Flow MR exam would impact the estimation of hemodynamic parameters, although some studies in other MRI contexts suggest that this may not be the case. Finally, the number of in-vivo datasets analyzed in this study is limited. These limitations motivate our future work, which will focus on three main areas. First, we will investigate using other purely viscous rheological models, such as Carreau-like, Cross, and Casson, for estimating hemodynamic parameters from velocity measurements obtained using 4D Flow MR imaging. Second, we will assess the impact of timely Hct measurements on estimating hemodynamic parameters from 4D Flow MR images. Finally, we will conduct a statistical analysis across a wide range of Hct values to evaluate the robustness and clinical applicability of the proposed methodology in both volunteers and patients.
@@ 7 Acknowledgments
The authors sincerely acknowledge the financial support provided by ANID through the following projects: Fondecyt de Iniciación en Investigación #11241098 (HM) and #11200481 (JS), Fondecyt Regular #1210156 (EC) and #1250287 (EC and FG), and the Millennium Science Initiative Program ICN2021_004 (JS). HM and JS also acknowledge the Department of Medical Imaging and Radiation Sciences at Monash University. FG also acknowledges the funding provided by the project DI VINCI PUCV 039.731/2025. The authors would like to express their sincere thanks to the Weierstraß-Institut für Angewandte Analysis und Stochastik (WIAS) for providing access to their servers for running CFD simulations.
@@ Appendix A Hemodynamic parameters for each aortic segments
Figure 1: Mean values and standard deviations of WSS, $\dot{E}_{L}$, and OSI within each aortic segment, estimated from simulated 4D Flow images generated from CFD simulations using power-law ($\mu_{\mathrm{PL}}$) and Newtonian fitted ($\mu_{\mathrm{NF}}$) viscosities. The estimations were performed using viscosity models and values consistent with those used in the CFD simulations that generated the synthetic 4D Flow images. Results are presented for both peak systole and diastole. Higher values and greater variability are observed in the ascending aorta due to the flow dynamics. The same differences shown in Figure 8 are clearly depicted here for each segment.